# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

### v4

@@ IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [7, 8] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [16].
@@ IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the semantic trajectory observation. The evaluator aims to identify deviations (if any) between the semantic trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [18].
@@ IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 3. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new semantic trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs “YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
@@ V-A Experimental Setup
In our experiment, the implementation of our proposed framework and comparison methods [5, 10, 7] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [20] as the foundational LLM. Based on the “o3-mini”, two separate LLMs are configured as the code generator and evaluator in our framework. The widely used UAV simulator, “AirSim” [27], is tailored to support the simulation component of our framework for code execution and state observations. The UAV model selected for experiment and evaluation is the default quadcopter in AirSim. The detailed implementation of our framework and experiment is available at our project website  [^1] Project Website: https://github.com/ai-uavsec/CLGSCE.
UAV Task Dataset: In our experiment, two categories of UAV tasks are selected to evaluate the performance of our framework under different operational complexities, i.e., Basic tasks and Advanced tasks. Each Basic task involves one or two intuitive actions to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). We use the Basic task dataset established by existing research [5], which contains 44 tasks in total. Differently, the Advanced tasks in our experiment require complex reasoning about the UAV’s position and orientation dynamics and typically require 6-19 actions to reach the goal state. We established the Advanced task dataset by emulating real-world UAV operation tasks, which consists of 20 tasks with different complexities. As summarized in Table I, there are four levels of complexity: “Low”, “Medium”, “High”, and “Very High”. The “Low” tasks involve flying simple geometric patterns. The “Medium” tasks introduce orientation requirements while flying the patterns (e.g., always orienting to the flying direction). The “High” tasks require flying complex patterns with orientation requirements (e.g., fly a figure 8 while orienting in the flying direction). The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flies the path reversely). The complete list of both Basic and Advanced tasks is available at our project website \@footnotemark All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
@@
Baseline Methods for Comparison: Three recently proposed methods [5, 10, 7] are selected as the baselines for comparison with our proposed approach, including
  • GSCE [5]: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement.
  • LLM Self-feedback (Self-Refine) [10]: A single LLM performs both code generation and self-refinement. The LLM analyzes semantic trajectory observations to detect deviations from the task and refines code based on its self-assessment.
  • Feedback with Numerical State Observations (Numerical) [7]: This method introduces a secondary LLM serving as evaluator designed according to section IV-D. The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement.
@@ V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table II. The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [28].
@@
Basic Task Advanced Task SR Completeness SR Completeness GSCE [5] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [10] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [7] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
For the Basic tasks, all methods showed remarkable performance. It is notable that the Numerical method [7] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness. This indicates that relying on numerical state observations can lead to misinterpretation in evaluation, even when handling simple tasks.
@@
Figure 6: Failure counts of different methods for each advanced task at different complexity levels
@@
TABLE III: Failure Cases Completeness in Each Level of Complexity
@@
GSCE [5] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
Self-Refine [10] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
Numerical [7] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
@@
Failure Case Analysis: To uncover the detailed performance of each method in its failed advanced tasks, we performed an in-depth failure case analysis. In particular, we aim to understand how many actions are correctly completed when a task fails. As summarized in Table III, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur. Differently, the average completeness of the compared baseline methods ranges from $48.5\%$ to $77.4\%$ for their failed tasks, which indicates errors occur for a significant portion of the required actions in these tasks. Fig. 6 further details the failure count on each failed task of all methods in our three rounds of experiment with task complexity and length of actions information. Our method has failures on only four tasks during three rounds of experiments. As a comparison, the baseline methods failed 11 (GSCE [5]), 17 (Self-Refine [10]), and 9 (Numerical [7]) tasks, respectively. In addition, Table III and Fig. 7 also show that our method has a stable high completeness of $90\%+$ for all failed tasks at different complexity levels. In contrast, the baseline methods have not only significantly lower average completeness ($21\%$ to $90.9\%$) on failed tasks at different levels but also a large variation in performance of different tasks.
@@
TABLE IV: Evaluator Configurations Description and Example System Prompts
@@ V-C2 Number of Refinement Iterations
In our proposed framework, the operation code generated by the Code Generator LLM will be refined through multiple iterations. The increased rounds of iterations provide additional opportunities to improve the code quality, however, they also introduce additional system cost (e.g., LLM API cost and processing time). To balance this tradeoff and identify an appropriate number of iterations, we performed experiments of our framework on advanced tasks with 0 to $10$ rounds of iterations. Our results presented in Fig. 8 show the changes of SR and completeness with the different number of iterations. Specifically, both SR and completeness are boosted after the iteration is introduced for refinement, and then have a slowly increasing trend until the $6\textsuperscript{th}$ iteration. After that, the performance of the subsequent iterations becomes relatively stable in terms of both SR and completeness, and even declines slightly. This is because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [29, 30], and thus may perform unnecessary changes after sufficient refinement has been performed. Our detailed examination identified that such a situation occurred on task 7 in our experiment. This observation demonstrates that iterative refinement can effectively contribute to the LLM-driven UAV operation code generation. Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff. According to our evaluation results, we set the maximum number of iterations to $6$. However, this number may need to be adjusted for different types of tasks. For example, setting a lower number when the tasks to complete have a relatively low complexity.
@@ V-D Evaluator Configuration Analysis
As the accuracy of the evaluation feedback directly influences the quality of code refinement, we analyze how the evaluator’s design influences both evaluation accuracy and overall system performance. We evaluated four configurations of the evaluator’s system prompt, including (1) roles, (2) roles ++ rules, (3) roles ++ references, and (4) roles ++ rules ++ references. As specified in Section IV-D, the roles define the LLM to be an evaluator and serve as the essential component to ensure the LLM will perform the evaluation task. Thus, we include the roles in each configuration of the evaluator. For each configuration, we instantiate the evaluator’s system prompt by including only the indicated components. The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table IV.
TABLE V: Performance over Evaluator Configurations

### v5

@@ IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [InteractivePlanning, CoPAL] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [NumberUnderstanding].
@@ IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the semantic trajectory observation. The evaluator aims to identify deviations (if any) between the semantic trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [LLM-bias].
@@ IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 3. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new semantic trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs “YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, human supervision should be involved to address potential errors in the code.
@@ V-A Experimental Setup
In our experiment, the implementation of our proposed framework and comparison methods [GSCE, SelfRefine, InteractivePlanning] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [o3mini] as the foundational LLM. Based on the “o3-mini”, two separate LLMs are configured as the code generator and evaluator in our framework. The widely used UAV simulator, “AirSim” [AirSim], is tailored to support the simulation component of our framework for code execution and state observations. The UAV model selected for experiment and evaluation is the default quadcopter in AirSim. The detailed implementation of our framework and experiment is available at our project website  [^1] Project Website: https://github.com/ai-uavsec/CLGSCE.
UAV Task Dataset: In our experiment, two categories of UAV tasks are selected to evaluate the performance of our framework under different operational complexities, i.e., Basic tasks and Advanced tasks. Each Basic task involves one or two intuitive actions to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). We use the Basic task dataset established by existing research [GSCE], which contains 44 tasks in total. Differently, the Advanced tasks in our experiment require complex reasoning about the UAV’s position and orientation dynamics and typically require 6-19 actions to reach the goal state. We established the Advanced task dataset by emulating real-world UAV operation tasks, which consists of 20 tasks with different complexities. As summarized in Table I, there are four levels of complexity: “Low”, “Medium”, “High”, and “Very High”. The “Low” tasks involve flying simple geometric patterns. The “Medium” tasks introduce orientation requirements while flying the patterns (e.g., always orienting to the flying direction). The “High” tasks require flying complex patterns with orientation requirements (e.g., fly a figure 8 while orienting in the flying direction). The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flying the path reversely). The complete list of both Basic and Advanced tasks is available at our project website 1. All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
@@
Baseline Methods for Comparison: Three recently proposed methods [GSCE, SelfRefine, InteractivePlanning] are selected as the baselines for comparison with our proposed approach, including
  • GSCE [GSCE]: An open-loop approach implemented by configuring a general LLM as UAV operation code generation agent (GSCE) with the system prompt that contains guidelines, skill APIs, constraints, and examples. The GSCE is first initialized with the system prompt; upon receiving a task prompt, it synthesizes control code based on the task instructions without any refinement.
  • LLM Self-feedback (Self-Refine) [SelfRefine]: A single LLM responsible for both code generation and refinement, which extends the GSCE method by self-evaluation and self-refinement of the code. Specifically, after the initial code is generated, it is executed in simulation to produce semantic trajectory observations, then the GSCE is prompted to evaluate these observations against the task description and iteratively refine the code based on its self-assessment.
  • Feedback with Numerical State Observations (Numerical) [InteractivePlanning]: Similar to our method, this method introduces a secondary LLM that evaluates the UAV states. In implementation, the code generator (GSCE) generates the initial code, which is then executed in simulation to produce UAV state observations in numerical representations of UAV state changes (format is the same as ground truth in the section V-B). Then, a separate LLM is configured as the evaluator with roles, rules, and references (numerical trajectory exemplars). The evaluator analyzes these numerical observations against the task description and provides corrective guidance to the code generator for subsequent refinement.
@@ V-B Evaluation Metrics
Tolerance: These metrics evaluate whether the generated code can correctly operate the UAV to perform the actions required by the task description. As AirSim simulation injects white noise in sensor measurements to emulate real-world sensing behavior [AirSim], there are offsets in the UAV state estimates produced by AirSim. To filter such noises that could affect trajectory matching with ground truth, we set a tolerance threshold of 1 meter for each positional parameter $[x,y,z]$ and 4 degrees for the Yaw $\theta$ in our implementation.
@@ V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table II. The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [LLMrandomness].
@@
Basic Task Advanced Task SR Completeness SR Completeness GSCE [GSCE] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [SelfRefine] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [InteractivePlanning] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
For the Basic tasks, all methods showed remarkable performance. It is notable that the Numerical method [InteractivePlanning] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness. This indicates that relying on numerical state observations can lead to misinterpretation in evaluation, even when handling simple tasks.
@@
TABLE III: Standard Deviation and $95\%$ Confidence Interval of Completeness for Tasks in Advanced Task Dataset
Task GSCE [GSCE] Self-Refine [SelfRefine] Numerical [InteractivePlanning] Ours
Standard Confidence Standard Confidence Standard Confidence Standard Confidence
Deviation Interval Deviation Interval Deviation Interval Deviation Interval
1 0 $[1.0,1.0]$ $0.385$ $[0.333,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
2 $0.128$ $[0.778,1.0]$ $0.449$ $[0.222,1.0]$ $0.421$ $[0.222,1.0]$ 0 $[1.0,1.0]$
3 0 $[1.0,1.0]$ $0.449$ $[0.222,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
4 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
5 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
6 $0.173$ $[0.700,1.0]$ $0.058$ $[0.9,1.0]$ $0.115$ $[0.8,1.0]$ 0 $[1.0,1.0]$
7 $0.167$ $[0.444,0.778]$ $0.321$ $[0.444,1.0]$ $0.280$ $[0.444,1.0]$ 0 $[1.0,1.0]$
8 $0.096$ $[0.778,0.944]$ $0.064$ $[0.889,1.0]$ $0.032$ $[0.944,1.0]$ $0.192$ $[0.611,0.944]$
9 $0.238$ $[0.588,1.0]$ $0.068$ $[0.412,0.529]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
10 $0.513$ $[0.111,1.0]$ 0 $[0.944,0.944]$ $0.289$ $[0.5,1.0]$ 0 $[1.0,1.0]$
11 0 $[1.0,1.0]$ 0 $[0.2,0.2]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
12 $0.052$ $[0.909,1.0]$ $0.472$ $[0.182,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
13 0 $[1.0,1.0]$ $0.527$ $[0.059,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
14 0 $[1.0,1.0]$ $0.064$ $[0.889,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
15 0 $[1.0,1.0]$ $0.289$ $[0.5,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
16 $0.102$ $[0.8,1.0]$ $0.346$ $[0.4,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
17 $0.058$ $[0.9,1.0]$ 0 $[0.4,0.4]$ $0.173$ $[0.6,0.9]$ $0.058$ $[0.9,1.0]$
18 $0.229$ $[0.579,1.0]$ 0 $[1.0,1.0]$ $0.169$ $[0.684,1.0]$ $0.030$ $[0.947,1.0]$
19 0 $[1.0,1.0]$ $0.272$ $[0.529,1.0]$ $0.238$ $[0.588,1.0]$ 0 $[1.0,1.0]$
20 0 $[0.944,0.944]$ 0 $[0.722,0.722]$ $0.032$ $[0.944,1.0]$ 0 $[0.944,0.944]$
Overall $0.164$ $[0.843,0.926]$ $0.325$ $[0.66,0.822]$ $0.178$ $[0.876,0.964]$ $\mathbf{0.055}$ $\mathbf{[0.969,0.995]}$
Statistical Analysis of Completeness: To evaluate the stability and consistency of our method, we performed statistical analysis for the completeness of advanced tasks. Specifically, the standard deviation and the $95\%$ confidence interval are measured for each advanced task across all three repetitions. As summarized in Table III, our method has a low overall standard deviation of $0.055$ and a high overall confidence interval $[0.969,0.995]$. These values also remain stable among almost all advanced tasks except for one of the failure cases. A more detailed analysis of the failure cases is also provided below. Compared with the baseline methods, which have higher standard deviations and wider confidence intervals for most tasks as shown in Table III, our method clearly achieves better stability and consistency.
Figure 6: Failure counts of different methods for each advanced task at different complexity levels.
@@
TABLE IV: Failure Cases Completeness in Each Level of Complexity
@@
GSCE [GSCE] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
Self-Refine [SelfRefine] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
Numerical [InteractivePlanning] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
@@
Failure Case Analysis: To uncover the detailed performance of each method in its failed advanced tasks, we performed an in-depth failure case analysis. In particular, we aim to understand how many actions are correctly completed when a task fails. As summarized in Table IV, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur. Differently, the average completeness of the compared baseline methods ranges from $48.5\%$ to $77.4\%$ for their failed tasks, which indicates errors occur for a significant portion of the required actions in these tasks. Fig. 6 further details the failure count on each failed task of all methods in our three rounds of experiment with task complexity and length of actions information. Our method has failures on only four tasks during three rounds of experiments. As a comparison, the baseline methods failed 11 (GSCE [GSCE]), 17 (Self-Refine [SelfRefine]), and 9 (Numerical [InteractivePlanning]) tasks, respectively. In addition, Table IV and Fig. 7 also show that our method has a stable high completeness of $90\%+$ for all failed tasks at different complexity levels. In contrast, the baseline methods have not only significantly lower average completeness ($21\%$ to $90.9\%$) on failed tasks at different levels but also a large variation in performance of different tasks.
@@
TABLE V: Evaluator Configurations Description and Example System Prompts
@@ V-C2 Number of Refinement Iterations
In our proposed framework, the operation code generated by the Code Generator LLM will be refined through multiple iterations. The increased rounds of iterations provide additional opportunities to improve the code quality. However, they also introduce additional system cost (e.g., LLM API cost and processing time). To balance this tradeoff and identify an appropriate number of iterations, we performed experiments of our framework on advanced tasks with 0 to $10$ rounds of iterations. As described in Algorithm 1, the early stopping is used in our design when the evaluator outputs “YES”, which prevents additional unnecessary refinement iterations.
The mean of our three repetition results is presented in Fig. 8, which shows the changes of SR and completeness with the increment of refinement iterations. Specifically, the mean iteration for our method to start converging with a relatively stable performance is 3. With a mean iteration of 6, our method is converged with the best performance in terms of both completeness and success rate. The performance declines slightly after the $6\textsuperscript{th}$ iteration because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [flipflop, LLM-behaviors] when receiving a negative query, and thus may perform unnecessary changes after sufficient refinement has been performed. This observation demonstrates that iterative refinement can effectively contribute to the LLM-driven UAV operation code generation. Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff between the success rate and system cost. As presented in Fig. 9, the computational cost and OpenAI API usage cost increase with the number of iterations. Thus, it is unnecessary to continue the refinement when the gain of success rate becomes very limited, i.e., after $6$ iterations in our setting. The number of iterations can be adjusted for different types of tasks. For example, setting a lower number when the tasks to complete have a relatively low complexity.
@@
Figure 9: Curves of per-task time and price over the number of refinement iterations.
@@ V-D Evaluator Configuration Analysis
As the accuracy of the evaluation feedback directly influences the quality of code refinement, we analyze how the evaluator’s design influences both evaluation accuracy and overall system performance. We evaluated four configurations of the evaluator’s system prompt, including (1) roles, (2) roles ++ rules, (3) roles ++ references, and (4) roles ++ rules ++ references. As specified in Section IV-D, the roles define the LLM to be an evaluator and serve as the essential component to ensure the LLM will perform the evaluation task. Thus, we include the roles in each configuration of the evaluator. For each configuration, we instantiate the evaluator’s system prompt by including only the indicated components. The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table V.
TABLE VI: Performance over Evaluator Configurations